### v1

# MedHal: An Evaluation Dataset for Medical Hallucination Detection

Abstract
We present MedHal, a novel large-scale dataset specifically designed to evaluate if models can detect hallucinations in medical texts. Current hallucination detection methods face significant limitations when applied to specialized domains like medicine, where they can have disastrous consequences. Existing medical datasets are either too small, containing only a few hundred samples, or focus on a single task like Question Answering or Natural Language Inference. MedHal addresses these gaps by: (1) incorporating diverse medical text sources and tasks; (2) providing a substantial volume of annotated samples suitable for training medical hallucination detection models; and (3) including explanations for factual inconsistencies to guide model learning. We demonstrate MedHal’s utility by training and evaluating a baseline medical hallucination detection model, showing improvements over general-purpose hallucination detection approaches. This resource enables more efficient evaluation of medical text generation systems while reducing reliance on costly expert review, potentially accelerating the development of medical AI research.
1 Introduction
Despite significant advancements in the field of Natural Language Processing, Large Language Models (LLMs) are prone to generating hallucinations (Huang et al. (2025)). This tendency poses a substantial obstacle to their application in multiple domains like medicine. While a robust hallucination detection is crucial, current hallucination metrics suffer from limitations in accuracy (Lin (2004); Banerjee & Lavie (2005); Zhang et al. (2019)). Traditional approaches relying on n-gram overlap (Lin (2004)) or entity overlap (Adams et al. (2024)) fail to capture the nuances of factual correctness. While more recent methods leverage general-purpose neural models (Kim et al. (2024); Laban et al. (2021)), incorporating context for potentially improved efficiency and have shown to correlate better with human judgment than n-gram overlap metrics (Zheng et al. (2023)), these models still exhibit suboptimal evaluation performance when applied to more specialized domain. Indeed, these models are mainly trained to detect general inconsistencies such as factual hallucinations or reasoning fallacies. In contexts like medical text generation, the most efficient way to detect hallucinated content is still to review generations by medical experts (Searle et al. (2023); Hegselmann et al. (2024)). However, this approach is costly and time-consuming, placing a significant burden on medical professionals. Therefore, a dedicated medical hallucination evaluation dataset is highly desirable. Such a resource could not only provide an accurate and relevant assessment of medical text generation, but also significantly reduce the cost and effort associated with medical AI research, facilitating its wider adoption. Current medical datasets often focus on single tasks like Question Answering (QA) or Natural Language Inference (NLI) to evaluate the hallucination levels of LLMs, limiting their scope and applicability to diverse medical text generation scenarios (Romanov & Shivade ; Pal et al. (2023); Chen et al. (2024); Yan et al. (2024); Agarwal et al. (2024); Hegselmann et al. (2024)). Furthermore, well-crafted medical hallucination datasets that currently exist only contain a few hundred samples, making them unsuitable for training large language models (LLMs) (Hegselmann et al. (2024)).
2 Related Work
Current state-of-the-art hallucination detection in the medical domain faces significant challenges due to the scarcity of comprehensive and diverse datasets. Existing hallucination datasets often focus on specific tasks or are limited in size. For instance, Med-Hallmark (Chen et al. (2024)) and Med-HVL (Yan et al. (2024)) focus on visual question-answering, assessing the ability of models to answer questions about medical images. MedHALU (Agarwal et al. (2024)) and Med-Halt (Pal et al. (2023)), on the other hand, are centered around QA and focus on evaluating LLMs’ ability to answer complex medical questions and detect inconsistencies in their responses. While the Hallucinations-MIMIC-DI dataset (Hegselmann et al. (2024)) provides a comprehensive annotation of discharge summaries and BHC sections, it only contains 100 samples, which is not sufficient to train a model or effectively evaluate another. Finally, some large datasets datasets like MedNLI exist for NLI tasks (Romanov & Shivade ). However, as shown in Figure 1, they rely on short, single-sentence premises and hypotheses, limiting their ability to capture complex medical reasoning and hallucinations.
Premise: Labs were notable for Cr 1.7 (baseline 0.5 per old records) and lactate 2.4.
Hypothesis: Patient has normal Cr
Figure 1: Sample from MedNLI
Our work differs from traditional hallucination medical datasets in several key aspects:
  1. 1.
    We utilize a greater variety of sources, such as clinical notes, clinical trials and medical questions, for the evaluation of hallucinations in more complex contexts
  2. 2.
    We create a large scale dataset that can be used to train a medical evaluator that can efficiently detect hallucinated content
  3. 3.
    We aim to provide explanations for why a statement is factual or not, a guiding signal that can be useful for LLM fine-tuning
3 Methodology
This study introduces MedHal, a novel dataset and benchmark designed for the evaluation of medical hallucination detection models. MedHal encompasses a diverse corpus of medical text, including clinical notes, research articles, and patient communication, annotated with various instances of factual inconsistencies and medical hallucinations. These hallucinations are generated through a suite of strategies tailored to individual task modalities, including answer substitution in question-answering (QA) and the introduction of conflicting statements in natural language inference (NLI). The efficacy of MedHal is demonstrated through the training of a baseline model and the comparative analysis of its performance against state-of-the-art models on the proposed benchmark. This resource aims to facilitate the development of more accurate and robust medical language models by providing a standardized and clinically relevant evaluation platform. Subsequent sections detail the construction of MedHal, describe the associated benchmark, and present experimental results demonstrating its utility in evaluating medical hallucination detection models. The methodological approach involves the transformation of existing medical datasets across diverse tasks (QA, Summarization, NLI, Information Extraction) into a unified hallucination detection task. This is achieved by framing the task as the binary classification of a given statement as factual or non-factual, potentially conditioned on a provided context.
3.1 Unified Task Formulation
To construct the dataset, a unified task formulation was established. Drawing inspiration from the NLI paradigm, each sample in the dataset is composed of a statement (analogous to the NLI hypothesis). This statement is classified as either factual or non-factual. The statement may be presented in the context of specific medical information (e.g., ”This patient has suffered from a myocardial infarction”) or as a general assertion within the medical domain (e.g., ”Aging causes an increase in blood pressure”). For each statement, the dataset provides a binary label indicating its factuality and, in the case of non-factual statements, an explanation detailing the specific inconsistency. A statement is defined as factual if all information contained within it is verifiable through the provided context or by established medical knowledge. Each sample in MedHal is structured as follows:
  • Statement: A proposition to be classified as factual or non-factual.
  • Context (optional): Contextual information relevant to the statement.
  • Factual (Binary): A binary label indicating the factuality of the statement (Yes/No).
  • Explanation: A textual explanation of the inconsistency for non-factual statements.
Table 1 shows examples of how statements are generated from samples.The following sections detail the generation of samples for this unified task, based on datasets from question-answering, information extraction, NLI, and summarization tasks.
Task Datasets Sample Generated Statement
Information Extraction Augmented-Clinical Notes A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis. She underwent arthroscopic surgery and was diagnosed with … $\rightarrow$ age: 10 years old The patient is 10 years old.
Summarization SumPubMed the large genotyping studies in the last decade have revolutionize genetic studies. our current ability to … $\rightarrow$ genetic admixture is a common caveat for genetic association analysis. these results… genetic admixture is a common caveat for genetic association analysis. these results…
NLI MedNLI Labs were notable for Cr 1.7 (baseline 0.5 per old records) and lactate 2.4. $\rightarrow$ Patient has normal Cr Patient has normal Cr
QA MedQA, MedMCQA Which of the following medications is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications? $\rightarrow$ Metformin Metformin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
Table 1: Example of samples are used to generate statements for each task
3.2 Question Answering Dataset Transformation
Question-answering (QA) datasets are structured around the presentation of a query followed by a set of potential responses, including binary (yes/no/maybe) and multiple-choice (A, B, C, …) formats. To generate factual samples from QA datasets, the question and its corresponding correct answer are transformed into a declarative statement using a large language model. Conversely, hallucinated samples are produced by pairing the question with incorrect answer options and subsequently converting these combinations into statements via the same LLM. Table 2 shows an example of a multiple-choice question converted into a statement. For generation, a consistent prompt template is employed for both factual and hallucinated sample generation. The prompt is given in Appendix 3. One-shot prompting is utilized to guide the LLM in accurately converting question-answer pairs into coherent statements. As for the explanation, certain QA datasets provide an explanation of why an answer is false. In these cases, we use the provided explanation for the non-factual statement. In other cases, we simply use the statement of the factual sample as the explanation.
Sample Type Question Answer Generated Statement
Factual Which of the following medications is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications? Metformin Metformin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
Non-Factual Insulin Insulin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
Table 2: Example of Question-Answering Dataset Transformation
3.3 Information Extraction Dataset Transformation
Information extraction (IE) datasets comprise, for each source document (e.g., clinical notes, clinical trials), a set of extracted text sequences related to specific concepts within that document. These extracted sequences, or ”extractions,” represent information pertaining to a particular attribute of the document. For example, from a clinical note, an extraction might represent the patient’s reason for the visit. To generate factual samples, the source document is used as the context, and the corresponding extraction is treated as the statement. Non-factual samples are generated by randomly interchanging extractions of the same value type between different documents (e.g., medications are swapped between patient records). When extractions are structured as key-value pairs, a large language model (LLM) is employed to transform these pairs into declarative statements. For instance, the key-value pair ”visit motivation: Lower back pain” is converted into the statement ”The patient’s visit motivation is lower back pain”. This transformation is performed using the prompt template shown in the Appendix 4 within a one-shot prompting framework. The explanation for a non-factual statement is provided by the factual extraction associated with the same document and key concept. An example of how a sample from an information extraction dataset is transformed is shown in Table 3. In cases where extractions are related to attributes with limited value diversity, such as sex, we ensure distinct values by explicitly switching them if random swapping results in identical values. The source documents are used without any preprocessing.
Sample Type Source Document Extraction Statement Explanation
Factual A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis… age: 10 years old The patient is 10 years old -
Non-Factual A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis… age: 16 years old The patient is 16 years old The patient is 10 years old
Table 3: Example of Information Extraction Dataset Transformation (the extraction from the non-factual statement is taken from another original sample)
3.4 Natural Language Inference Dataset Transformation
Given the inherent similarity between the proposed task and natural language inference (NLI), NLI datasets are directly adaptable to the MedHal benchmark. We utilize NLI datasets specifically related to the medical domain. Factual samples are derived from hypotheses in entailment examples, while hallucinated samples are generated from hypotheses in contradiction examples. The premise component of the NLI example serves as the context for the corresponding statement. Samples resulting in a neutral label are ignored during dataset construction. The NLI datasets are not preprocessed beyond this filtering. Note that in this case, we do not have explanations for the non-factual samples.
3.5 Summarization Dataset Transformation
For summarization datasets, factual samples are constructed using genuine summaries or sentences extracted from them. Hallucinated samples are generated by extracting a single sentence from the original summary and employing an LLM to modify it, introducing self-contradictory information. The modified, hallucinated sentence is then reintegrated into the summary at its original position. The source text intended for summarization serves as the context, and the summary itself is treated as the statement. The explanation for a non-factual statement is simply the original sentence used to generate the hallucinated version. Sentences for hallucination generation are selected randomly, with a minimum length of 100 characters to ensure sufficient contextual information for the LLM. An example of how a sample is generated is shown in Table 4. The prompt used to generate self-contradictory sentences is shown in Appendix 5.The goal of this process is to evaluate if models can effectively detect small mistakes in statements consisting of long sequences of text.
Sample Type Source Document Summary Statement Explanation
Factual a central feature in the maturation of hearing is a transition in the electrical signature of cochlear hair cells from spontaneous calcium… cochlear hair cells are high-frequency sensory receptors… cochlear hair cells are high-frequency sensory receptors… -
Non-Factual cochlear hair cells are low-frequency sensory receptors… According to the source document, cochlear hair cells are high-frequency sensory receptors…
Table 4: Example of Summarization Dataset Transformation
3.6 Dataset Description
To ensure a balanced benchmark, each factual sample is paired with a corresponding non-factual sample. Furthermore, the benchmark is constructed using a diverse set of datasets, encompassing all task modalities previously described, to ensure variability in document sources. Table 5 presents a comprehensive overview of the datasets utilized in the generation of this benchmark.
Dataset Task S Content Type Source # Samples # Gen
MedMCQA QA Medical Content Pal et al. (2022) 183,000 70,730
MedNLI NLI Clinical Notes Herlihy & Rudinger (2021) 11,232 7,488
ACM IE Clinical Notes Bonnet & Boulenger (2024) 22,000 73,040
MedQA QA Medical Content Jin et al. (2020) 12,723 18,906
PubMedSum Sum Clinical Trials Gupta et al. (2021) 33,772 178,657
Table 5: Description of datasets used to generate the MedHal benchmark (S is for Synthetic)
The discrepancy between the original sample count and the generated sample count is attributable to several factors. Firstly, for certain tasks like information extraction, original samples can yield multiple generated samples. Secondly, to manage computational resource limitations, the total number of generated samples was reduced. Finally, samples resulting in token sequences longer than 8192 were excluded to mitigate potential adverse effects on model training. Preprocessing was selectively applied to specific datasets. For the ACM dataset, only valid JSON-structured extractions were retained. For QA datasets, explanations containing phrases such as ”Answer is c)” that referred to the original dataset options were removed using regular expressions, as we are using these samples outside the QA setting. For instance, the explanation ”Answer is c), Metformin is most commonly used as a first-line treatement, because…)” would be transformed to ”Metformin is most commonly used as a first-line treatement, because…)”.
The MedHal benchmark is partitioned into training, validation, and test sets according to an 80/10/10 ratio, ensuring a consistent evaluation framework. This results in 313,920 samples for the training set, 17458 samples in the validation set and 17443 samples in the test set.
4 Experiments
4.1 Models
For the generation of the MedHal dataset, we utilize the Llama-3-70B model (Grattafiori et al. (2024)), a state-of-the-art language model known for its strong performance across various language generation tasks. For all prompting strategies employed in the sample generation process, a one-shot learning approach was adopted to provide the model with examples of the desired output format prior to generation. Importantly, to ensure a fair evaluation of our proposed benchmark, the dataset generation was performed exclusively using the training splits of the individual source datasets. This approach allows for the independent evaluation of our trained models on the original test sets of the constituent datasets.
4.2 Evaluator
To demonstrate the utility of the MedHal dataset, we fine-tuned a dedicated model for the detection of medical hallucinated content. This evaluator model is initialized with the Llama-1B model (Grattafiori et al. (2024)) and subsequently fine-tuned on the training set of the MedHal dataset. The detailed training configuration is provided in Appendix 7. The selection of a 1 billion parameter model as the base for our evaluator is motivated by the desire to provide a resource that is accessible to a wider range of researchers and domain experts, including those with limited computational resources. Existing evaluator models and general-purpose language models often comprise several billion parameters, necessitating significant infrastructure for their deployment. This can restrict the practical application of LLM-based evaluation metrics in certain research contexts. Furthermore, the use of a smaller model contributes to lower infrastructure costs associated with training and deployment. Future work will explore the potential for enhancing the evaluation metrics through the utilization of larger models or the optimization of the fine-tuning process.
4.3 Results
We also fine-tuned a general-purpose Llama-3 1B parameter model (MedHal-Llama-1B) on the dataset. This serves as a baseline to evaluate whether specialized medical LLMs, which have undergone extensive medical domain fine-tuning, already demonstrate strong performance in detecting hallucinated content and can effectively function as judges in medical hallucination detection tasks. The prompt used for all non fine-tuned models is shown in Appendix 6. We report two types of metrics: factuality metrics and explanation metrics. Factuality metrics aim to measure if the model correctly detected hallucinated content. These include the precision, recall, and F1-score. Explanation metrics aim to evaluate if explanations generated by models when statements are not factual are valid. They evaluate if models, when detecting hallucinated content, correctly identify which part of the statement is hallucinated. In order to compute these metrics, we only take into account samples where both the prediction and the ground truth indicate a non-factual statements since in other cases, a true explanation might not exist or not be generated. Explanation metrics include the ROUGE-1 (R1) and ROUGE-2 (R2) (Lin (2004)) scores and the BLEU (Papineni et al. (2002)) score. Results are shown in Table 6.
FT Type Models Factuality Explanation
P R F1 BLEU R1 R2
General Llama-1B-Instruct 0.53 0.32 0.40 0.01 0.13 0.03
Llama-8B-Instruct 0.59 0.62 0.60 0.05 0.24 0.14
Medical BioMistral-7B 0.56 0.43 0.49 0.03 0.22 0.11
MedLlama-8B 0.52 0.59 0.55 0.03 0.21 0.08
OpenBioLLM-8B 0.52 0.77 0.62 0.04 0.21 0.10
General MedHal-Llama (Ours) 0.75 0.77 0.76 0.45 0.70 0.59
Table 6: Results of models on test set of MedHal (FT: Fine-Tuned, P: Precision, R: Recall, F1: F1-Score, R1: ROUGE-1, R2: ROUGE-2)
As models do not always follow the output format given in the prompt in Appendix 6, we only compute the factuality metrics on samples that matched the output format. We also compute the F1-score of each model, based on the origin of each sample to see if models tend to excel in certain tasks compared to other tasks. These results are shown in Figure 2.
Figure 2: Accuracy of models based on each dataset source
5 Discussion
5.1 Overall Model Performance
Our evaluation reveals several key insights regarding the performance of various models in detecting and explaining medical hallucinations. Notably, general-purpose models demonstrated a surprising aptitude for identifying factual inaccuracies, with Llama-8B-Instruct achieving the highest precision (0.59) in models that were not fine-tuned. However, part of this performance might also be due to the memorization of samples available in our dataset as it is built from datasets training sets. We base this hypothesis on the observation that this model significantly outperformed others by a wide margin on MedNLI, the only dataset that had not been transformed by another LLM. In contrast, specialized medical models, despite extensive fine-tuning on medical texts, exhibited lower performance in factuality metrics. Only OpenBioLLM demonstrated higher overall performance with an F1-score of 0.62. However, it exhibited a higher recall than precision, indicating a tendency towards overconfidence in identifying factual statements. Interestingly, general-purpose models also outperformed medical models in generating explanations for identified hallucinations, as evidenced by higher BLEU and ROUGE scores. After manual evaluation, this is also mainly due to medical models not generally following instructions to explain their choices.
5.2 Task-Specific Performance Variations
Performance varied across different tasks, highlighting the challenges associated with specific data sources. Notably, information extraction proved challenging for all models, including our fine-tuned MedHal-Llama-1B, which performed only slightly better than random chance. We hypothesize that this might be attributable to the synthetic nature of the information extraction dataset, which might introduce some inaccuracies in our labels. Furthermore, we observed no significant performance disparity between tasks that require contextual analysis (information extraction, summarization, NLI) and tasks that operate on isolated statements (question-answering). This suggests that the presence or absence of contextual information does not substantially impact the models’ ability to detect and explain hallucinations.
5.3 Fine-Tuned Evaluation Model
Fine-tuning a 1 billion parameter Llama model (MedHal-Llama) on our dataset yielded impressive results, surpassing the performance of models seven to eight times its size, including those with extensive medical fine-tuning. This finding might indicate that specialized medical fine-tuning does not guarantee a higher performance in hallucination detection, at least as measured by our dataset. MedHal-Llama especially demonstrated high performance in detecting hallucinations within the summarization task in both factuality and explanation metrics. This is likely due to the overrepresentation of summarization samples in our dataset, leading the model to correctly identifying which sentences were falsified. Future research should however focus on balancing the dataset across all tasks to ensure a more equitable evaluation.
6 Limitations and Future Work
It is important to acknowledge some limitations of our methodology, the main one being that we rely on generated content. The validity of the generated hallucinated content was primarily assessed through visual inspection of a sample of the generated data and relied on the performance of a 70 billion model at not hallucinating over small sequences of texts. Indeed, all generations only involved generating a small number of tokens for the model. However, ideally, a comprehensive evaluation involving medical professionals would be conducted to validate the accuracy and clinical relevance of the generated dataset. Additionally, while we only fine-tune a general purpose model, future work should also evaluate the performance of fine-tuning medical models to this task to see how they compare to our current fine-tuned model that is based on a general purpose model.
7 Conclusion
In this study, we introduced MedHal, the first large-scale dataset designed to specifically address the challenge of hallucination detection across diverse clinical tasks, including summarization, natural language inference, information extraction, and question answering. We evaluated a range of medical and general-purpose models on this dataset, providing a comprehensive benchmark for assessing their performance in this critical domain. Furthermore, we fine-tuned an evaluator model on MedHal, which we make available to the community as a baseline and tool for future research. While MedHal represents a significant step forward in addressing the limitations of existing hallucination detection methods in the medical field, we acknowledge the need for further refinement. Specifically, future work should prioritize validating the accuracy and clinical relevance of labels and explanations through expert reviews. Additionally, it should also focus on balancing the dataset across all included tasks to ensure a more equitable evaluation. By addressing these limitations, we can enhance the reliability and utility of MedHal.
8 Acknowledgements
We gratefully acknowledge the financial support provided by the FRQS (Fonds de Recherche du Québec – Santé) Chair in Artificial Intelligence and Digital Health. This research also benefited from the computational resources and GPU infrastructure provided by the Digital Research Alliance of Canada (DRAC) and Calcul Québec. We extend our sincere thanks to Prof. John Kildea for his valuable input during initial discussions on the challenges of working in clinical settings.
Appendix A Appendix
system: Given a question and an answer, your role is to transform the question into a statement by incorporating the answer with it. Do not add any details that is not mentioned in the question or the answer. user: Question: Which of the following agents is most commonly associated with recurrent meningitis due to CSF leaks? Answer: Pneumococci assistant: Pneumococci is most commonly associated with recurrent meningitis due to CSF leaks user: Question: [question] Answer: [answer]
Figure 3: Prompt template used for the QA task
system: You are tasked with transforming structured medical data into natural language statements about a patient. Each input will contain 4 elements: - concept: The type of information being described (e.g., dosage, age, symptoms) - value: The specific information or measurement - category: The broad medical category this information belongs to (e.g., treatment, patient information, symptoms) - concept_reference: The specific element that the value refers to (e.g., a specific medication, a specific symptom) Your task is to generate a clear, grammatically correct sentence that conveys this information in a medical context. Follow these rules: 1. Use appropriate verbs based on the concept: - For treatments: ’takes’, ’receives’, ’is prescribed’ - For symptoms: ’experiences’, ’reports’, ’presents with’ - For measurements/states: ’is’, ’has’, ’shows’ - For time-related concepts: ’has been’, ’started’, ’continues’ 2. Incorporate the concept_reference when it adds clarity 3. Use present tense 4. Maintain medical terminology as provided 5. When the concept_reference is ’None’ or does not add clarity, don’t include it in the statement 6. The statement should be a single sentence. Do not include any other information in the statement aside from the concept and the extraction. Only output the statement and nothing else. user: category: patient medical history value: History of left elbow arthrodesis performed for posttraumatic arthritis at the age of 18, concept_reference: None assistant: The patient underwent left elbow arthrodesis as a treatment for posttraumatic arthritis when they were 18 years old. user: category: [category] value: [value] concept_reference: [concept_reference]
Figure 4: Prompt template used for the IE task (for more information, see Bonnet & Boulenger (2024))
user: You will be given a text and a sentence that was extracted from the text. Your task is to transform the sentence by introducing a deliberate inaccuracy. Strategies can include: - Changing numerical values - Inverting the meaning - Using antonyms - Negating the original statement Text: [text] Sentence: [sentence] Ensure the new sentence remains grammatically correct but semantically different from the original. Only output the transformed sentence without any additional text.
Figure 5: Prompt template used for the summarization task
### Task Description - You will evaluate whether a medical statement is factually accurate. - The statement may reference a provided context. - Respond with ”YES” if the statement is factually correct or ”NO” if it contains inaccuracies. - In order to answer YES, everything in the statement must be supported by the context. - In order to answer NO, there must be at least one piece of information in the statement that is not supported by the context. - You must also provide an explanation of why you think the statement is factual or not. If it is factual, put ”The statement is factual” as your explanation. - Your answer should follow the following format : Factual: [YES/NO] Explanation: [Your explanation] ### Context [context] ### Statement [statement]
Figure 6: Prompt template used for evaluating models
load_in_4bit: true max_seq_len: 8192 per_device_train_batch_size: 8 per_device_eval_batch_size: 8 num_train_epochs: 1 gradient_accumulation_steps: 2 optim: paged_adamw_8bit r: 16 lora_alpha: 16
Figure 7: Training configuration with QLora

### v2

@@
 Abstract
-We present MedHal, a novel large-scale dataset specifically designed to evaluate if models can detect hallucinations in medical texts.
+We present MedHal [^1] MedHal is distributed under Apache 2.0 license.
+Link will be provided uppon acceptance., a novel large-scale dataset specifically designed to evaluate if models can detect hallucinations in medical texts.
 Current hallucination detection methods face significant limitations when applied to specialized domains like medicine, where they can have disastrous consequences.
-Existing medical datasets are either too small, containing only a few hundred samples, or focus on a single task like Question Answering or Natural Language Inference.
+Existing medical datasets generally do not focus on hallucination detection and are either too small, containing only a few hundred samples, or focus on a single task like Question Answering or Natural Language Inference.
 MedHal addresses these gaps by: (1) incorporating diverse medical text sources and tasks; (2) providing a substantial volume of annotated samples suitable for training medical hallucination detection models; and (3) including explanations for factual inconsistencies to guide model learning.
@@
 This resource enables more efficient evaluation of medical text generation systems while reducing reliance on costly expert review, potentially accelerating the development of medical AI research.
+MedHal: An Evaluation Dataset for Medical Hallucination Detection
+Gaya Mehenni1,2,3, Fabrice Lamarche1,2,3, Odette Rios-Ibacache 4,5, John Kildea4,5, Amal Zouaq1,2,3 1LAMA-WeST Lab, 2Polytechnique Montréal 3Mila - Quebec AI Institute 4Medical Physics Unit - RI-MUHC 5McGill University Correspondence: gaya.mehenni@polymtl.ca
 1 Introduction
-Despite significant advancements in the field of Natural Language Processing, Large Language Models (LLMs) are prone to generating hallucinations (Huang et al.
-(2025)).
+Despite significant advancements in the field of Natural Language Processing, Large Language Models (LLMs) are prone to generating hallucinations Huang et al.
+(2025).
 This tendency poses a substantial obstacle to their application in multiple domains like medicine.
-While a robust hallucination detection is crucial, current hallucination metrics suffer from limitations in accuracy (Lin (2004); Banerjee & Lavie (2005); Zhang et al.
-(2019)).
-Traditional approaches relying on n-gram overlap (Lin (2004)) or entity overlap (Adams et al.
-(2024)) fail to capture the nuances of factual correctness.
-While more recent methods leverage general-purpose neural models (Kim et al.
+While robust hallucination detection is essential, existing hallucination metrics continue to suffer from accuracy limitations Lin (2004); Banerjee and Lavie (2005); Zhang et al.
+(2019).
+Traditional methods based on n-gram overlap Lin (2004) or entity overlap Adams et al.
+(2024) often fail to capture the nuances of factual correctness.
+Although more recent approaches employ general-purpose neural models Kim et al.
 (2024); Laban et al.
-(2021)), incorporating context for potentially improved efficiency and have shown to correlate better with human judgment than n-gram overlap metrics (Zheng et al.
-(2023)), these models still exhibit suboptimal evaluation performance when applied to more specialized domain.
-Indeed, these models are mainly trained to detect general inconsistencies such as factual hallucinations or reasoning fallacies.
-In contexts like medical text generation, the most efficient way to detect hallucinated content is still to review generations by medical experts (Searle et al.
+(2021) and leverage contextual information for potentially improved evaluation, showing stronger correlation with human judgment compared to n-gram-based metrics Zheng et al.
+(2023), their performance remains suboptimal in specialized domains.
+These models are primarily trained to detect broad inconsistencies, such as factual inaccuracies or reasoning errors.
+In high-stakes contexts like medical text generation, the most reliable method for identifying hallucinated content remains expert human review Searle et al.
 (2023); Hegselmann et al.
-(2024)).
-However, this approach is costly and time-consuming, placing a significant burden on medical professionals.
-Therefore, a dedicated medical hallucination evaluation dataset is highly desirable.
-Such a resource could not only provide an accurate and relevant assessment of medical text generation, but also significantly reduce the cost and effort associated with medical AI research, facilitating its wider adoption.
-Current medical datasets often focus on single tasks like Question Answering (QA) or Natural Language Inference (NLI) to evaluate the hallucination levels of LLMs, limiting their scope and applicability to diverse medical text generation scenarios (Romanov & Shivade ; Pal et al.
-(2023); Chen et al.
+(2024b).
+However, this process is both time-consuming and resource-intensive, placing a considerable burden on medical professionals.
+Consequently, there is a clear need for a dedicated medical hallucination evaluation dataset—both to assess the performance of LLMs on hallucination detection tasks and to enable fine-tuning for improved hallucination detection capabilities.
+Currently, many medical datasets focus on narrow tasks such as Question Answering (QA) or Natural Language Inference (NLI), limiting their utility across the full range of medical text generation scenarios Romanov and Shivade ; Chen et al.
 (2024); Yan et al.
-(2024); Agarwal et al.
 (2024); Hegselmann et al.
-(2024)).
-Furthermore, well-crafted medical hallucination datasets that currently exist only contain a few hundred samples, making them unsuitable for training large language models (LLMs) (Hegselmann et al.
-(2024)).
+(2024b).
+Moreover, existing domain-specific hallucination datasets typically contain only a few hundred samples, rendering them insufficient for training or robustly evaluating large language models (LLMs) Hegselmann et al.
+(2024b).
+In this work, we propose MedHal, a dataset for the accurate and domain-relevant assessment of generated medical content, significantly reducing the cost and effort required for evaluation—ultimately facilitating broader adoption of medical AI systems.
 2 Related Work
@@
 Existing hallucination datasets often focus on specific tasks or are limited in size.
-For instance, Med-Hallmark (Chen et al.
-(2024)) and Med-HVL (Yan et al.
-(2024)) focus on visual question-answering, assessing the ability of models to answer questions about medical images.
-MedHALU (Agarwal et al.
-(2024)) and Med-Halt (Pal et al.
-(2023)), on the other hand, are centered around QA and focus on evaluating LLMs’ ability to answer complex medical questions and detect inconsistencies in their responses.
-While the Hallucinations-MIMIC-DI dataset (Hegselmann et al.
-(2024)) provides a comprehensive annotation of discharge summaries and BHC sections, it only contains 100 samples, which is not sufficient to train a model or effectively evaluate another.
-Finally, some large datasets datasets like MedNLI exist for NLI tasks (Romanov & Shivade ).
+For instance, Med-Hallmark Chen et al.
+(2024) and Med-HVL Yan et al.
+(2024) focus on visual question-answering, assessing the ability of models to answer questions about medical images.
+MedHaluAgarwal et al.
+(2024) and Med-Halt Pal et al.
+(2023), on the other hand, are centered around QA and focus on evaluating LLMs’ ability to answer complex medical questions and detect inconsistencies in their responses.
+While the Hallucinations-MIMIC-DI dataset Hegselmann et al.
+(2024b) provides a comprehensive annotation of discharge summaries and BHC sections, it only contains 100 samples, which is not sufficient to train a model or effectively evaluate another.
+Finally, some large datasets like MedNLI exist for NLI tasks Romanov and Shivade .
 However, as shown in Figure 1, they rely on short, single-sentence premises and hypotheses, limiting their ability to capture complex medical reasoning and hallucinations.
@@
 This statement is classified as either factual or non-factual.
-The statement may be presented in the context of specific medical information (e.g., ”This patient has suffered from a myocardial infarction”) or as a general assertion within the medical domain (e.g., ”Aging causes an increase in blood pressure”).
+The statement may be presented in the context of specific medical information (e.g., "This patient has suffered from a myocardial infarction") or as a general assertion within the medical domain (e.g., "Aging causes an increase in blood pressure").
 For each statement, the dataset provides a binary label indicating its factuality and, in the case of non-factual statements, an explanation detailing the specific inconsistency.
@@
 $\rightarrow$ Metformin Metformin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
-Table 1: Example of samples are used to generate statements for each task
+Table 1: Example of samples used to generate statements for each task
 3.2 Question Answering Dataset Transformation
-Question-answering (QA) datasets are structured around the presentation of a query followed by a set of potential responses, including binary (yes/no/maybe) and multiple-choice (A, B, C, …) formats.
+Question-answering (QA) datasets are structured around a query followed by a set of potential responses, including binary (yes/no/maybe) and multiple-choice (A, B, C, …) formats.
 To generate factual samples from QA datasets, the question and its corresponding correct answer are transformed into a declarative statement using a large language model.
 Conversely, hallucinated samples are produced by pairing the question with incorrect answer options and subsequently converting these combinations into statements via the same LLM.
-Table 2 shows an example of a multiple-choice question converted into a statement.
+Table 11 in Appendix A shows an example of a multiple-choice question converted into a statement.
 For generation, a consistent prompt template is employed for both factual and hallucinated sample generation.
@@
 In other cases, we simply use the statement of the factual sample as the explanation.
-Sample Type Question Answer Generated Statement
-Factual Which of the following medications is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications?
-Metformin Metformin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
-Non-Factual Insulin Insulin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
-Table 2: Example of Question-Answering Dataset Transformation
 3.3 Information Extraction Dataset Transformation
 Information extraction (IE) datasets comprise, for each source document (e.g., clinical notes, clinical trials), a set of extracted text sequences related to specific concepts within that document.
-These extracted sequences, or ”extractions,” represent information pertaining to a particular attribute of the document.
+These extracted sequences, or "extractions," represent information pertaining to a particular attribute of the document.
 For example, from a clinical note, an extraction might represent the patient’s reason for the visit.
@@
 When extractions are structured as key-value pairs, a large language model (LLM) is employed to transform these pairs into declarative statements.
-For instance, the key-value pair ”visit motivation: Lower back pain” is converted into the statement ”The patient’s visit motivation is lower back pain”.
+For instance, the key-value pair "visit motivation: Lower back pain" is converted into the statement "The patient’s visit motivation is lower back pain".
 This transformation is performed using the prompt template shown in the Appendix 3 within a one-shot prompting framework.
 The explanation for a non-factual statement is provided by the factual extraction associated with the same document and key concept.
-An example of how a sample from an information extraction dataset is transformed is shown in Table 3.
+An example of how a sample from an information extraction dataset is transformed is shown in Table 12 in Appendix A.
 In cases where extractions are related to attributes with limited value diversity, such as sex, we ensure distinct values by explicitly switching them if random swapping results in identical values.
 The source documents are used without any preprocessing.
-Sample Type Source Document Extraction Statement Explanation
-Factual A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis… age: 10 years old The patient is 10 years old -
-Non-Factual A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis… age: 16 years old The patient is 16 years old The patient is 10 years old
-Table 3: Example of Information Extraction Dataset Transformation (the extraction from the non-factual statement is taken from another original sample)
 3.4 Natural Language Inference Dataset Transformation
@@
 Sentences for hallucination generation are selected randomly, with a minimum length of 100 characters to ensure sufficient contextual information for the LLM.
-An example of how a sample is generated is shown in Table 4.
-The prompt used to generate self-contradictory sentences is shown in Appendix 5.The goal of this process is to evaluate if models can effectively detect small mistakes in statements consisting of long sequences of text.
+An example of how a sample is generated is shown in Table 13 in Appendix A.
+The prompt used to generate self-contradictory sentences is shown in Appendix 4.
+The goal of this process is to evaluate if models can effectively detect fine-grained hallucinations in statements consisting of long sequences of text.
+Dataset Task S Content Type Source # Samples # Gen
+MedMCQA QA Medical Content Pal et al. (2022) 183,000 278,021
+MedNLI NLI Clinical Notes Herlihy and Rudinger (2021) 11,232 7,488
+ACM IE Clinical Notes Bonnet and Boulenger (2024) 22,000 327,568
+MedQA QA Medical Content Jin et al. (2020) 12,723 18,680
+PubMedSum Sum Clinical Trials Gupta et al. (2021) 33,772 195,339
+Table 2: Description of datasets, after filtering, used to create the MedHal benchmark (S is for Synthetic)
+3.6 Model
+For the creation of the MedHal dataset, we utilize the Llama-3-70B model Grattafiori et al.
+(2024), a state-of-the-art language model known for its strong performance across various language generation tasks.
+For all prompting strategies employed in the sample generation process, a one-shot learning approach was adopted to provide the model with examples of the desired output format prior to generation.
+See Sections 3.2, 3.3, 3.4, and 3.5 for details on what is generated for each task.
+Importantly, to ensure a fair evaluation of our proposed benchmark, the dataset creation was performed exclusively using the training splits of the individual source datasets.
+This approach allows for the independent evaluation of our trained models on the original test sets of the constituent datasets.
+3.7 Data Quality Assessment
+To ensure data quality, we incorporated a filtering step where a Llama-3.3-70B model assessed the factuality of each generated statement, given its context.
+For tasks with subtle nuances, such as QA and summarization, the model was provided with the correct answer or the original sentence, respectively, to guide its judgment.
+For NLI tasks, it simply evaluated the statement’s factuality.
+Samples were retained if the model’s assessment matched the dataset’s assigned label; otherwise, they were removed.
+Table 3 presents the percentage of samples retained for each dataset after this filtering process.
+A majority of samples were retained across most datasets, with the exception of ACM, which is the only synthetic dataset in our collection.
+Dataset % Samples Kept
+SumPubMed 99.30
+ACM 82.00
+MedMCQA 96.87
+MedQA 99.44
+MedNLI 100.00
+Table 3: Percentage of samples kept after the filtering step for each dataset.
+3.8 MedHal Dataset Description
+To ensure a balanced benchmark, each factual sample is paired with a corresponding non-factual sample.
+Furthermore, the benchmark is constructed using a diverse set of datasets, encompassing all task modalities previously described, to ensure variability in document sources.
+Table 2 presents a comprehensive overview of the datasets utilized in the generation of this benchmark.
+The MedHal benchmark is distributed in several configurations to accommodate diverse research needs.
+The full dataset comprises a total of 827,096 samples.
+Additionally, we provide a length-filtered (LF) version, retaining only samples where the combined length of the context and statement is less than 30,000 characters.
+This specific length constraint is chosen to align with the 8,192 token context length common to most current state-of-the-art 1-10B parameter models, reflecting our study’s focus on the performance of smaller language models in medical settings.
+Furthermore, we release a balanced version (BAL) of the dataset in terms of the tasks.
+In this configuration, the Question Answering (QA), Summarization, and Information Extraction (IE) tasks each contain approximately 150,000 samples.
+The NLI task, however, could not be balanced to this extent due to the inherently limited number of available samples for that specific task.
+Table 4 details the number of samples and their respective proportions for each task across the different dataset versions.
+Task Full LF LF + BAL
+Summarization 0.24 0.19 0.31
+Information Extraction 0.40 0.42 0.31
+Question Answering 0.36 0.38 0.35
+NLI 0.01 0.01 0.02
+Total samples 827,096 779,376 469,025
+Table 4: Sample proportions for each task across different dataset versions (LF=Length Filtered, BAL=Balanced)
+3.9 Expert Validation
+To validate the quality of the filtered dataset, one of the authors (PhD student in Medical Physics), conducted a manual annotation of a subset of the dataset.
+We randomly selected 50 examples from each Task (25 for each Factual and Non-Factual labels [^2] We only select Non-Factual samples for the Summarization task since Factual Samples contain no modification of the data.).
+The annotator is then tasked with validating the label, considering both the Context and Statement.
+A sample is considered valid if its label is judged as correct by the annotator.
+Annotation guidelines are presented in Appendix D.
+Statistics describing the validity of each task is presented in Table 5.
+We observe near perfect validity for all tasks, and manual observation of the invalid examples show samples that lack relevance or important pieces of information from the context.
+Task Valid samples (%)
+Question Answering 0.91
+Information Extraction 0.96
+Natural Language Inference 0.98
+Summarization 1.0
+Table 5: Expert validation results per task.
+Type Model Factuality Explanation
+P R F1 BLEU R1 R2 N
+General Llama-3.2-1B 0.51 0.41 0.46 0.01 0.08 0.03 1040
+Llama-3-8B 0.66 0.58 0.62 0.03 0.12 0.08 15962
+Medical BioMistral-7B 0.64 0.51 0.57 0.09 0.25 0.19 13365
+MedLlama-8B 0.94 0.26 0.40 0.09 0.32 0.22 22395
+OpenBioLLM-8B 0.61 0.76 0.68 0.01 0.04 0.01 7153
+Evaluator Prometheus-2-8x7B 0.85 0.69 0.76 - - - -
+HallOumi-8B 0.67 0.76 0.72 - - - -
+Table 6: Performance of models on MedHal’s test set (N refers to the number of samples where an explanation was extracted when prompting the model)
+4 Experiments
+4.1 General Evaluation
+To determine how well current AI models detect medical hallucinations, we evaluated their performance on the MedHal test set.
+This assessment helps us identify which models are most adept at recognizing hallucinated content within a medical context and, more importantly, provides insights into whether specific fine-tuning strategies improve medical hallucination detection.
+We compare the performance of general-purpose models (Llama-3.2-1B, Llama-3-8B), medical models fine-tuned on general medical datasets (BioMistral-7B, MedLlama-8B, OpenBioLLM-8B), and evaluator models fine-tuned to judge answers of other LLMs (Proemtheus-2-8x7B, HallOumi-8B) .
+For general-purpose and medical models, we used the prompt template detailed in Figure 5.
+For Prometheus-2-8x7B and HallOumi-8B, we adhered to the prompt formats recommended by their original authors, as these are optimized for their fine-tuned performance.
+We use two main types of metrics to evaluate the models: factuality metrics and explanation metrics.
+Factuality metrics measure how accurately a model identifies factual versus non-factual content.
+This includes common measures like precision, recall, and F1-score.
+Due to inconsistencies in model output when using the prompt format from Figure 5.
+You can see these results in Table 6.
+Explanation metrics assess the validity of the explanations that the models provide for non-factual statements.
+Specifically, these metrics check if a model, after identifying non-factual content, correctly pinpoints the exact erroneous part of the statement.
+The explanation metrics are ROUGE-1 (R1), ROUGE-2 (R2) Lin (2004), and BLEU Papineni et al.
+(2002) scores.
+To ensure a valid comparison, we only consider samples where both the model’s prediction and the ground truth label indicate a non-factual statement.
+This guarantees that a true explanation exists and that the model attempted to generate one.
+4.2 Impact of Specialized Training
+We fine-tuned several models on the MedHal training set to assess the impact of specialized training on their performance.
+Our main goal was to see if fine-tuning a medical model specifically for medical hallucination detection, or specializing a general hallucination detection model on medical data, would yield better results.
+We also investigated whether simply fine-tuning a general-purpose model could achieve similar performance, potentially bypassing the need for more specialized initial training.
+For this experiment, we fine-tuned Llama-3-8B, Llama-3-OpenBioLLM-8B, and HallOumi-8B on MedHal.
+The prompt used during training is shown in Figure 6.
+All models were fine-tuned using the same QLora configuration [^3] For details on training setup, see Appendix C.
+We report the same metrics detailed in Table 6, along with the difference in F1 score between the fine-tuned and non-fine-tuned versions of each model in Table 7.
+[…]
+P R F1 BLEU R1 R2
+Llama-3-8B 0.93 0.84 0.88 0.52 0.81 0.74 +0.26
+OpenBioLLM-8B 0.90 0.88 0.90 0.53 0.81 0.74 +0.22
+HallOumi-8B 0.92 0.80 0.86 0.51 0.80 0.72 +0.14
+[…]
+4.3 Task Ablation
+We perform an ablation on the task training data to better understand the individual contribution of each task to the models’ overall performance.
+In this setting, we fine-tune the Llama-3-8B base model separately on the training samples for each task: Summarization, Question Answering, and Information Extraction.
+The same prompt format is used across all experiments to ensure consistency.
+We exclude the Natural Language Inference (NLI) task from this analysis because it contains significantly fewer samples compared to the others, which would make a direct comparison of its impact inequitable.
+By isolating each task, this ablation allows us to assess its relative importance.
+A substantial performance gain from a single task would indicate its high value, whereas a decrease or negligible improvement compared to the baseline could suggest the task is less critical for the overall goal.
+Results are shown in Table 8.
+[…]
+P R F1 BLEU R1 R2
+Summarization 0.71 0.95 0.81 0.46 0.69 0.62 +0.19
+Question Answering 0.78 0.67 0.72 0.19 0.50 0.38 +0.10
+Information Extraction 0.73 0.72 0.74 0.17 0.54 0.41 +0.12
+[…]
+4.4 Downstream Performance
+We assess the validity of our MedHal dataset and its utility for benchmarking by evaluating the downstream task performance of models fine-tuned on MedHal on MedNLI Herlihy and Rudinger (2021) and another hallucination detection dataset Hegselmann et al.
+(2024a).
+4.4.1 MedNLI
+We first evaluate base model on MedNLI and compare them against a Llama-3-8B model fine-tuned on MedHal.
+Additionally, we establish a baseline by fine-tuning a Llama-3-8B model on the MedNLI training set for 5 epochs to validate the impact of our tasks on performance.
+As models fine-tuned on our dataset cannot detect neutral statements in MedNLI, we filter out these samples from the test set during evaluation of all models.
+We report the F1 scores of all models in Table 9.
+Fine-Tuning Base Model F1-Score
+None Llama-3.2-1B 0.64
+Llama-3-8B 0.65
+BioMistral-7B 0.56
+MedLlama-3-8B 0.66
+OpenBioLLM-8B 0.64
+Prometheus-2-8x7B 0.62
+HallOumi-8B 0.89
+MedNLI Llama-3-8B 0.95
+MedHal Llama-3-8B 0.96
+Table 9: F1-Score of models on the MedNLI test set (samples with a neutral labels have been filtered out)
+4.4.2 Hallucination Dataset
+This benchmark Hegselmann et al.
+(2024a) is derived from MIMIC-III and contains summaries of patient notes generated by frontier models, which were then annotated for factual inconsistencies by medical students.
+We consider a summary as non-factual if it contains any of the eleven types of factual inconsistencies in the dataset, including ’Fact Contradicted’, ’Unsupported Medication’, or ’Unsupported Procedure’.
+Although the dataset contains only 210 samples—152 of which include hallucinations—its recent release makes it highly unlikely that the evaluated models were exposed to it during training.
+We compare 1-shot prompting strategy using several models - as models consistently failed to identify hallucinations in a zero-shot setting-, a base Llama-3-8B model fine-tuned on the MedHal train set to measure the impact of our dataset, and evaluation models (Prometheus-2-8x7B, HallOumi-8B).
+Results are shown in Table 10.
+Model Precision Recall F1
+Llama-3.2-1B (1 shot) 0.81 0.32 0.45
+Llama-3.1-8B (1 shot) 1.00 0.13 0.22
+OpenBioLLM-8B (1 shot) 0.85 0.60 0.70
+Prometheus-2-8x7B 0.77 0.75 0.76
+HallOumi-8B 0.91 0.61 0.73
+MedHal (Llama-3-8B) 0.95 0.63 0.76
+Table 10: Evaluation of different models on an Hallucination Dataset
+5 Discussion
+5.1 Performance of Models
+Our evaluation reveals several key insights into the performance of various models in detecting and explaining medical hallucinations.
+A notable finding is the clear superiority of models specifically fine-tuned as evaluators in the task of factuality detection.
+Prometheus-2-8x7B emerges as the most accurate model, achieving the highest F1-score of 0.76, followed closely by HallOumi-8B with an F1-score of 0.72.
+Surprisingly, general-purpose models demonstrate a strong aptitude for this task, with Llama-3-8B (F1=0.62) outperforming some specialized medical models.
+The performance of medical models is mixed; while OpenBioLLM-8B surpasses the general-purpose baseline with an F1-score of 0.68, other models in this category lag.
+When assessing the quality of explanations for non-factual content, however, the trend reverses.
+Here, medical models show a distinct advantage, with MedLlama-8B significantly outperforming all other models across explanation metrics (R1=0.32, R2=0.22) while having the highest number of valid explanations (indicating a strong capacity to follow instructions).
+We perform a detailed analysis of model errors in Appendix E.
+5.2 Task-Specific Performance Variations
+Table 8 shows that each task in the dataset contributes uniquely to improving model performance in detecting factual and non-factual content, with all tasks boosting the baseline F1 score by at least 10%.
+Among them, the Summarization task shows the most significant impact.
+Fine-tuning solely on Summarization leads to substantially greater gains over the baseline than fine-tuning on Question Answering or Information Extraction.
+These improvements are consistent across both "Factuality" and "Explanation" metrics.
+The effectiveness of Summarization can be attributed to its inherent complexity.
+Unlike Question Answering, which typically draws on general medical knowledge, Summarization requires a more complex understanding of the context and the statement.
+While Information Extraction also involves contextual analysis, it deals with short, focused statements.
+Summarization, in contrast, involves longer statements and minor inconsistencies, requiring the model to analyze both the statement and the broader context.
+This might lead to the model learning more robust, generalizable representations for hallucination detection, which also benefit tasks like Information Extraction and some Question Answering cases.
+Representations learned from Question Answering and Information Extraction do not transfer as effectively to other tasks in the dataset.
+5.3 Evaluation on Downstream Datasets
+The results on the MedNLI dataset, shown in Table 9, highlight the benefits of MedHal fine-tuning.
+Our Llama-3-8B model achieves an F1-score of 0.96, a 7% improvement over the best baseline (HallOumi-8B), and slightly outperforms the Llama-3-8B model fine-tuned solely on MedNLI (F1=0.95), despite only encountering MedNLI samples once during multi-task training, compared to 5 epoch for Llama-3-8B.
+This small margin may be due to performance saturation on MedNLI for models of this size, but it still underscores the value of diverse task signals.
+On the MIMIC-based hallucination dataset, our MedHal-tuned model achieves an F1 score of 0.76, matching Prometheus-2-8x7B despite being nearly six times smaller.
+While Prometheus balances precision (0.77) and recall (0.75), our model shows a high-precision profile, achieving 0.95 precision and 0.63 recall.
+In clinical settings, such high precision is highly valuable, ensuring flagged statements are rarely false positives.
+In contrast, models like Llama-3.1-8B (1-shot) achieve perfect precision but suffer from extremely low recall (0.13), limiting practical use.
+Our model also surpasses others like OpenBioLLM-3-8B (F1=0.70) and HallOumi-8B (F1=0.73), reaffirming the strength of MedHal’s multi-task training approach.
+Conclusion
+In this study, we introduced MedHal, the first large-scale dataset designed to specifically address the challenge of hallucination detection across diverse clinical tasks, including summarization, natural language inference, information extraction, and question answering.
+We evaluated a range of medical and general-purpose models on this dataset, providing a comprehensive benchmark for assessing their performance in this critical domain.
+Furthermore, we fine-tuned multiple models on MedHal to evaluate the impact of specialized pre-training on medical hallucination detection performance.
+Plus, we also assess which tasks models tend to struggle more with.
+MedHal represents a significant step forward in addressing the limitations of existing hallucination detection methods in the medical field.
+Limitations
+It is important to acknowledge some limitations of our methodology, the main one being that we rely on generated content.
+Ideally, a more comprehensive evaluation involving medical professionals could be conducted to validate the accuracy and clinical relevance of the generated dataset.
+We also note that while our expert validation showed great data quality, a multi-annotator setup would allow measurement of agreement metrics and would improve our validity claims.
+Plus, we could incorporate more complex multi-hop reasoning hallucination detection tasks into the dataset.
+This would allow us to identify error patterns and find more room for improvement in developing hallucination detection models.
+Finally, this resource is only available in English and augmenting it with other languages would be beneficial.
+This is left for future work.
+Appendix
+Appendix A Data Transformation Examples
+Sample Type Question Answer Generated Statement
+Factual Which of the following medications is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications?
+Metformin Metformin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
+Non-Factual Insulin Insulin is most commonly used as first-line treatment for newly diagnosed type 2 diabetes mellitus in patients without contraindications.
+Table 11: Example of Question-Answering Dataset Transformation
+Sample Type Source Document Extraction Statement Explanation
+Factual A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis… age: 10 years old The patient is 10 years old -
+Non-Factual A 10-year-old girl first noted a swollen left knee and underwent repeated arthrocentesis… age: 16 years old The patient is 16 years old The patient is 10 years old
+Table 12: Example of Information Extraction Dataset Transformation (the extraction from the non-factual statement is taken from another original sample)
 Sample Type Source Document Summary Statement Explanation
@@
 Table 13: Example of Summarization Dataset Transformation
-3.6 Dataset Description
-To ensure a balanced benchmark, each factual sample is paired with a corresponding non-factual sample.
-Furthermore, the benchmark is constructed using a diverse set of datasets, encompassing all task modalities previously described, to ensure variability in document sources.
-Table 5 presents a comprehensive overview of the datasets utilized in the generation of this benchmark.
-Dataset Task S Content Type Source # Samples # Gen
-MedMCQA QA Medical Content Pal et al. (2022) 183,000 70,730
-MedNLI NLI Clinical Notes Herlihy & Rudinger (2021) 11,232 7,488
-ACM IE Clinical Notes Bonnet & Boulenger (2024) 22,000 73,040
-MedQA QA Medical Content Jin et al. (2020) 12,723 18,906
-PubMedSum Sum Clinical Trials Gupta et al. (2021) 33,772 178,657
-Table 5: Description of datasets used to generate the MedHal benchmark (S is for Synthetic)
-The discrepancy between the original sample count and the generated sample count is attributable to several factors.
-Firstly, for certain tasks like information extraction, original samples can yield multiple generated samples.
-Secondly, to manage computational resource limitations, the total number of generated samples was reduced.
-Finally, samples resulting in token sequences longer than 8192 were excluded to mitigate potential adverse effects on model training.
-Preprocessing was selectively applied to specific datasets.
-For the ACM dataset, only valid JSON-structured extractions were retained.
-For QA datasets, explanations containing phrases such as ”Answer is c)” that referred to the original dataset options were removed using regular expressions, as we are using these samples outside the QA setting.
-For instance, the explanation ”Answer is c), Metformin is most commonly used as a first-line treatement, because…)” would be transformed to ”Metformin is most commonly used as a first-line treatement, because…)”.
-The MedHal benchmark is partitioned into training, validation, and test sets according to an 80/10/10 ratio, ensuring a consistent evaluation framework.
-This results in 313,920 samples for the training set, 17458 samples in the validation set and 17443 samples in the test set.
-4 Experiments
-4.1 Models
-For the generation of the MedHal dataset, we utilize the Llama-3-70B model (Grattafiori et al.
-(2024)), a state-of-the-art language model known for its strong performance across various language generation tasks.
-For all prompting strategies employed in the sample generation process, a one-shot learning approach was adopted to provide the model with examples of the desired output format prior to generation.
-Importantly, to ensure a fair evaluation of our proposed benchmark, the dataset generation was performed exclusively using the training splits of the individual source datasets.
-This approach allows for the independent evaluation of our trained models on the original test sets of the constituent datasets.
-4.2 Evaluator
-To demonstrate the utility of the MedHal dataset, we fine-tuned a dedicated model for the detection of medical hallucinated content.
-This evaluator model is initialized with the Llama-1B model (Grattafiori et al.
-(2024)) and subsequently fine-tuned on the training set of the MedHal dataset.
-The detailed training configuration is provided in Appendix 7.
-The selection of a 1 billion parameter model as the base for our evaluator is motivated by the desire to provide a resource that is accessible to a wider range of researchers and domain experts, including those with limited computational resources.
-Existing evaluator models and general-purpose language models often comprise several billion parameters, necessitating significant infrastructure for their deployment.
-This can restrict the practical application of LLM-based evaluation metrics in certain research contexts.
-Furthermore, the use of a smaller model contributes to lower infrastructure costs associated with training and deployment.
-Future work will explore the potential for enhancing the evaluation metrics through the utilization of larger models or the optimization of the fine-tuning process.
-4.3 Results
-We also fine-tuned a general-purpose Llama-3 1B parameter model (MedHal-Llama-1B) on the dataset.
-This serves as a baseline to evaluate whether specialized medical LLMs, which have undergone extensive medical domain fine-tuning, already demonstrate strong performance in detecting hallucinated content and can effectively function as judges in medical hallucination detection tasks.
-The prompt used for all non fine-tuned models is shown in Appendix 6.
-We report two types of metrics: factuality metrics and explanation metrics.
-Factuality metrics aim to measure if the model correctly detected hallucinated content.
-These include the precision, recall, and F1-score.
-Explanation metrics aim to evaluate if explanations generated by models when statements are not factual are valid.
-They evaluate if models, when detecting hallucinated content, correctly identify which part of the statement is hallucinated.
-In order to compute these metrics, we only take into account samples where both the prediction and the ground truth indicate a non-factual statements since in other cases, a true explanation might not exist or not be generated.
-Explanation metrics include the ROUGE-1 (R1) and ROUGE-2 (R2) (Lin (2004)) scores and the BLEU (Papineni et al.
-(2002)) score.
-Results are shown in Table 6.
-FT Type Models Factuality Explanation
-P R F1 BLEU R1 R2
-General Llama-1B-Instruct 0.53 0.32 0.40 0.01 0.13 0.03
-Llama-8B-Instruct 0.59 0.62 0.60 0.05 0.24 0.14
-Medical BioMistral-7B 0.56 0.43 0.49 0.03 0.22 0.11
-MedLlama-8B 0.52 0.59 0.55 0.03 0.21 0.08
-OpenBioLLM-8B 0.52 0.77 0.62 0.04 0.21 0.10
-General MedHal-Llama (Ours) 0.75 0.77 0.76 0.45 0.70 0.59
-Table 6: Results of models on test set of MedHal (FT: Fine-Tuned, P: Precision, R: Recall, F1: F1-Score, R1: ROUGE-1, R2: ROUGE-2)
-As models do not always follow the output format given in the prompt in Appendix 6, we only compute the factuality metrics on samples that matched the output format.
-We also compute the F1-score of each model, based on the origin of each sample to see if models tend to excel in certain tasks compared to other tasks.
-These results are shown in Figure 2.
-Figure 2: Accuracy of models based on each dataset source
-5 Discussion
-5.1 Overall Model Performance
-Our evaluation reveals several key insights regarding the performance of various models in detecting and explaining medical hallucinations.
-Notably, general-purpose models demonstrated a surprising aptitude for identifying factual inaccuracies, with Llama-8B-Instruct achieving the highest precision (0.59) in models that were not fine-tuned.
-However, part of this performance might also be due to the memorization of samples available in our dataset as it is built from datasets training sets.
-We base this hypothesis on the observation that this model significantly outperformed others by a wide margin on MedNLI, the only dataset that had not been transformed by another LLM.
-In contrast, specialized medical models, despite extensive fine-tuning on medical texts, exhibited lower performance in factuality metrics.
-Only OpenBioLLM demonstrated higher overall performance with an F1-score of 0.62.
-However, it exhibited a higher recall than precision, indicating a tendency towards overconfidence in identifying factual statements.
-Interestingly, general-purpose models also outperformed medical models in generating explanations for identified hallucinations, as evidenced by higher BLEU and ROUGE scores.
-After manual evaluation, this is also mainly due to medical models not generally following instructions to explain their choices.
-5.2 Task-Specific Performance Variations
-Performance varied across different tasks, highlighting the challenges associated with specific data sources.
-Notably, information extraction proved challenging for all models, including our fine-tuned MedHal-Llama-1B, which performed only slightly better than random chance.
-We hypothesize that this might be attributable to the synthetic nature of the information extraction dataset, which might introduce some inaccuracies in our labels.
-Furthermore, we observed no significant performance disparity between tasks that require contextual analysis (information extraction, summarization, NLI) and tasks that operate on isolated statements (question-answering).
-This suggests that the presence or absence of contextual information does not substantially impact the models’ ability to detect and explain hallucinations.
-5.3 Fine-Tuned Evaluation Model
-Fine-tuning a 1 billion parameter Llama model (MedHal-Llama) on our dataset yielded impressive results, surpassing the performance of models seven to eight times its size, including those with extensive medical fine-tuning.
-This finding might indicate that specialized medical fine-tuning does not guarantee a higher performance in hallucination detection, at least as measured by our dataset.
-MedHal-Llama especially demonstrated high performance in detecting hallucinations within the summarization task in both factuality and explanation metrics.
-This is likely due to the overrepresentation of summarization samples in our dataset, leading the model to correctly identifying which sentences were falsified.
-Future research should however focus on balancing the dataset across all tasks to ensure a more equitable evaluation.
-6 Limitations and Future Work
-It is important to acknowledge some limitations of our methodology, the main one being that we rely on generated content.
-The validity of the generated hallucinated content was primarily assessed through visual inspection of a sample of the generated data and relied on the performance of a 70 billion model at not hallucinating over small sequences of texts.
-Indeed, all generations only involved generating a small number of tokens for the model.
-However, ideally, a comprehensive evaluation involving medical professionals would be conducted to validate the accuracy and clinical relevance of the generated dataset.
-Additionally, while we only fine-tune a general purpose model, future work should also evaluate the performance of fine-tuning medical models to this task to see how they compare to our current fine-tuned model that is based on a general purpose model.
-7 Conclusion
-In this study, we introduced MedHal, the first large-scale dataset designed to specifically address the challenge of hallucination detection across diverse clinical tasks, including summarization, natural language inference, information extraction, and question answering.
-We evaluated a range of medical and general-purpose models on this dataset, providing a comprehensive benchmark for assessing their performance in this critical domain.
-Furthermore, we fine-tuned an evaluator model on MedHal, which we make available to the community as a baseline and tool for future research.
-While MedHal represents a significant step forward in addressing the limitations of existing hallucination detection methods in the medical field, we acknowledge the need for further refinement.
-Specifically, future work should prioritize validating the accuracy and clinical relevance of labels and explanations through expert reviews.
-Additionally, it should also focus on balancing the dataset across all included tasks to ensure a more equitable evaluation.
-By addressing these limitations, we can enhance the reliability and utility of MedHal.
-8 Acknowledgements
-We gratefully acknowledge the financial support provided by the FRQS (Fonds de Recherche du Québec – Santé) Chair in Artificial Intelligence and Digital Health.
-This research also benefited from the computational resources and GPU infrastructure provided by the Digital Research Alliance of Canada (DRAC) and Calcul Québec.
-We extend our sincere thanks to Prof.
-John Kildea for his valuable input during initial discussions on the challenges of working in clinical settings.
-Appendix A Appendix
+Appendix B Prompt templates
 system: Given a question and an answer, your role is to transform the question into a statement by incorporating the answer with it.
@@
 user: category: [category] value: [value] concept_reference: [concept_reference]
-Figure 4: Prompt template used for the IE task (for more information, see Bonnet & Boulenger (2024))
+Figure 3: Prompt template used for the IE task (for more information, see Bonnet and Boulenger (2024))
 user: You will be given a text and a sentence that was extracted from the text.
@@
 - The statement may reference a provided context.
-- Respond with ”YES” if the statement is factually correct or ”NO” if it contains inaccuracies.
+- Respond with "YES" if the statement is factually correct or "NO" if it contains inaccuracies.
 - In order to answer YES, everything in the statement must be supported by the context.
@@
 - You must also provide an explanation of why you think the statement is factual or not.
-If it is factual, put ”The statement is factual” as your explanation.
+If it is factual, put "The statement is factual" as your explanation.
 - Your answer should follow the following format : Factual: [YES/NO] Explanation: [Your explanation] ### Context [context] ### Statement [statement]
 Figure 5: Prompt template used for evaluating models
+### Task Description - You will evaluate whether a medical statement is factually accurate.
+- The statement may reference a provided context.
+- Respond with "YES" if the statement is factually correct or "NO" if it contains inaccuracies.
+- In order to answer YES, everything in the statement must be supported by the context.
+- In order to answer NO, there must be at least one piece of information in the statement that is not supported by the context.
+- You must also provide an explanation of why you think the statement is factual or not.
+If it is factual, put "The statement is factual" as your explanation.
+### Context [context] ### Statement [statement]
+Figure 6: Prompt template used for training models
+Appendix C Training configuration
 load_in_4bit: true max_seq_len: 8192 per_device_train_batch_size: 8 per_device_eval_batch_size: 8 num_train_epochs: 1 gradient_accumulation_steps: 2 optim: paged_adamw_8bit r: 16 lora_alpha: 16
 Figure 7: Training configuration with QLora
+Appendix D Annotation Guidelines
+Evaluation Criteria
+What Constitutes a “Factual” Statement?
+A statement is considered factual if and only if:
+  1. 1.
+    Every piece of information mentioned in the statement can be directly supported by the provided context, OR
+  2. 2.
+    When no context is provided, the statement represents well-established, accurate medical knowledge.
+  3. 3.
+    There are no contradictions between the statement and the context.
+  4. 4.
+    The statement does not include unsupported claims or extrapolations beyond what the context states.
+Evaluation Process
+  1. 1.
+    Review the Context
+    Carefully read the provided context (e.g., research abstract, clinical information, etc.)
+  2. 2.
+    Analyze the Statement
+    Break down the statement into individual claims and check each claim against the context.
+  3. 3.
+    Cross-Reference with Medical Knowledge
+    For statements without context, evaluate them against established medical facts.
+  4. 4.
+    Make Your Assessment
+    Fill in the evaluation columns:
+    • Valid Column:
+      • YES: You agree with the current label (TRUE for factual / FALSE for non-factual)
+      • NO: You disagree with the current label
+    • Comment Column: Provide your reasoning, including:
+      • Specific discrepancies found (if any)
+      • Medical knowledge that supports or contradicts the statement
+      • Concerns about accuracy or completeness
+Example Evaluation
+Context:
+“A randomized controlled trial of 200 patients found that Drug A reduced hospital readmissions by 30% compared to placebo (p<0.05p<0.05).”
+Statement:
+“Drug A significantly reduces hospital readmissions by 30% in all patients.”
+Current Label: TRUE
+Your Evaluation:
+  • Valid: NO
+  • Comment: “While the 30% reduction is accurate, the statement overgeneralizes by claiming effectiveness ‘in all patients’ when the study only included 200 patients with specific characteristics.
+The word ‘significantly’ is supported by p<0.05p<0.05.”
+Appendix E Error Analysis
+To uncover the underlying reasons why models fail to identify factual and non-factual statements, we perform an error analysis by categorizing each incorrect factuality prediction by its original task.
+Figure 8: Proportion of Model Errors by Task Source
+The results shown in Figure 8 reveal a consistent pattern: the vast majority of errors for most models originate from samples derived from the Question Answering task.
+This trend likely originates from the nature of these samples, which often assess general medical knowledge rather than requiring contextual analysis of a provided document.
+This suggests that while the models may possess the necessary reasoning capabilities to detect inconsistencies within a given text, they frequently lack the specific, ingrained medical knowledge required to validate standalone factual assertions.
+Suprisingly, this knowledge-versus-reasoning gap is mainly illustrated by OpenBioLLM-8B, a medical model that performs well on contextual tasks but whose errors are almost exclusively from the QA task.
+This profile suggests that its failures are not due to flawed logic but to a deficit in its medical knowledge.
+In contrast, a smaller model like Llama-3.2-1B displays a more uniform error distribution across all tasks, indicating more fundamental limitations in both its reasoning and knowledge faculties.
+Ultimately, this analysis indicates that a primary obstacle for many current small models is not a failure of reasoning itself, but insufficient medical knowledge, a weakness that is most exposed when contextual clues are absent and pure factual recall is demanded.
+Additionally, we also investigate whether medical hallucination detection is mainly linked to context length.
+To do so, we plot the distribution of context lengths across samples that were not correctly labelled by models.
+Figure 9: Distribution of context lengths across samples that were not correctly labelled by models in the test set
+Figure 9 illustrates the frequency distribution of context lengths for samples where models did not correctly label the sample, relative to the overall distribution of context lengths in the test set.
+Several key trends emerge from this visualization.
+Firstly, OpenBioLLM seems to make mistakes on samples with shorter context lengths.
+This finding aligns with observations from Figure 8, as the Question Answering task typically involves shorter contexts compared to others.
+Second, several models (BioMistral-7B, Llama-3-8B, MedLlama-8B, Prometheus-2-8x7B, HallOumi-8B) have a higher peek density at the first mode than the second mode contrary to the test set distribution.
+This trend is particularly surprising, as models generally tend to struggle more with longer contexts.
+These results indicate that the task is a better indicator of the errors of a model than the actual context length.
+Only Llama-3.2-1B seems to have a distribution similar to the test set’s, indicating that context length does not have a big impact on performance.
+This might be due to current models supporting a larger context length.
+Indeed, our samples might not test to the limit the context retrieval capabilities of models.